OpenStudy (anonymous):

if you see -12-14 +8-8+9-10+24_&= , how do you know if the signs are for signed numbers or are they operational signs ?

OpenStudy (anonymous):

I am not sure what you mean... I see each as operational signs since there are no multiplications or ( ).. . I am assuming 24_& is your mistake????

OpenStudy (anonymous):

Think of it as combining rather than adding and subtracting. Combining two negatives yields a negative, combing two positives yields a positives. combing a negative and a positive yields whichever is greater. When you see something like -1-(-1) it can be rewritten as -1+1.

OpenStudy (anonymous):

that is a good way to look at it.

OpenStudy (anonymous):

thanks you're right 24_& was my mistake. i understand -1-(-1)=-1+1 and i see it this way almost always,but wouldn't this be right? -1 is a signed number and - or more appropriatly (-) be the indicated operation therefore -1(-)-1=-1(+)+1=o be more appropriate, more right and less confusing and the intention far more clear. if memory serves me right this is the way it used to be done untill some math school book writer decided to add another confusing shortcut and wonder why our math understanding is droping.